OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown.

OpenStudy (anonymous):

y = two fifthsx + 1 y = -five halvesx - twenty-seven halves y = -two fifthsx - 3 y = five halvesx + twenty-three halves

OpenStudy (anonymous):

since u have a figure... you can determine the equation of the given line.... some points are visible already... standard form of line equation... \[Ax + By + C = 0\]

OpenStudy (anonymous):

in slope-intercept form: \[y = -\frac{ A }{ B }x - \frac{ C }{ B }\]

OpenStudy (anonymous):

the perpendicular line will have a standard form... \[Bx + Ay + D = 0\]

OpenStudy (anonymous):

the slope-intercept form.... \[y = -\frac{ B }{ A }x - \frac{ D }{ A }\]

OpenStudy (anonymous):

... and this line contains the point mentioned in the problem...

OpenStudy (anonymous):

... that will lead u to right answer...

OpenStudy (anonymous):

i am so confused. :$

OpenStudy (anonymous):

ok let us get the equation of the given line... obvious points are (0,2), (2,7), (-3,-1)...

OpenStudy (anonymous):

to check, solve for the slope is constant at chosen obvious points... let P1(2,7), P2(0,2), P3(-3,-1)... \[m = \frac{ (2 - 7) }{ (0 - 2) } = \frac{ -5 }{ -2 } = \frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

another 2 points to make sure... \[m = \frac{ (-3 - 2) }{ (-2 - 0) } = \frac{ -5 }{ -2 } = \frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

correction for P3, it should be (-2,-3)... not (-3,-1)

OpenStudy (anonymous):

so for the given line, the slope \[m = \frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

using slope-intercept form... \[y = \frac{ 5 }{ 2 }x + 2\]

OpenStudy (anonymous):

in standard form... \[5x - 2y + 4 = 0\]

OpenStudy (agent0smith):

Okay first, do you know how to find the slope of that line? you can easily find it from the graph... remember rise over run?

OpenStudy (anonymous):

our A = 5, B = -2 and C = 4

OpenStudy (agent0smith):

@Orion1213 slow down and stop confusing her with unnecessary stuff :P

OpenStudy (anonymous):

For perpendicular line, we interchange A and B (property of perpendicular line)... A = -2, B = 5, therefore in standard form... \[-2x + 5y + D = 0\]

OpenStudy (anonymous):

what is D?

OpenStudy (anonymous):

@almartinez26 kindly react if u have a question.... :)

OpenStudy (agent0smith):

@almartinez26

OpenStudy (anonymous):

sorry i lost my connection for a while...

OpenStudy (anonymous):

D is the constant of the perpendicular line... we can solve by considering the given point (-5,-1) wherein the perpendicular line passed thru...

OpenStudy (anonymous):

It's all good. Thanks for your help but agent0smith explained it to me, i got it now.

OpenStudy (anonymous):

ok... no need to continue i think... :)